A 12.40 gram sample is heated to an initial temperature of 50 oC. The sample is then placed in a calorimeter containing 6 grams of water starting at 24 oC. If the final temperature is 35 oC what is the specific heat of the sample?

sH2O = 4.18


1
Expert's answer
2021-03-15T09:05:31-0400

Solution:

1) This problem can be summarized thusly:

qlost by sample = qgained by water

 

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J oC-1 g-1)

Tf = final temperature (oC)

Ti = initial temperature (oC)


2) Therefore:

qsample = (12.40) × Csample × (35 - 50) = -186 × Csample

qwater = (6.00 g) × (4.18) × (35 - 24) = 275.88

 

275.88 = 186 × Csample

Csample = 1.483 J oC-1 g-1


Answer: The specific heat of the sample is 1.483 J oC-1 g-1.